OpenStudy (anonymous):

x^(2)10^x-2x(10^x)=3(10^x) solve equation..

OpenStudy (anonymous):

10^x(x^2 -2x -3) =0 x=1,-3

OpenStudy (anonymous):

can you walk me through it please

OpenStudy (anonymous):

that was wrong

OpenStudy (anonymous):

x^(2)10^x-2x(10^x)=3(10^x)

OpenStudy (anonymous):

Factor out 10^x then solve like any other quadratic equation

OpenStudy (anonymous):

so 10^x(x^2-2x)=3(10^x)?

OpenStudy (anonymous):

Bring LHS and RHS to same side and solve quadratin x^2 -2x -3

OpenStudy (anonymous):

bring what?

OpenStudy (anonymous):

all terms to one side

OpenStudy (anonymous):

so 10^x(x^2-2x-3)=0

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yep now you have it factorize quadratic and 10^x gives no roots as it is always greater than zero

OpenStudy (anonymous):

so final answer is -1,3

OpenStudy (anonymous):

1,-3

OpenStudy (anonymous):

no -1,3 ...why u get 1,-3?

OpenStudy (anonymous):

YEah it was a test